# Transitions to Intermittent Chaos in Quorum Sensing Dynamics

### v2

@@ Abstract
This study analyses the dynamical consequences of heterogeneous temporal delays within a quorum sensing-inspired (QS-inspired) system, specifically addressing the differential response kinetics of two subpopulations to signalling molecules. A nonlinear delay differential equation (DDE) model, predicated upon an activator-inhibitor framework, is formulated to represent the interspecies interactions. Key analytical techniques, including the derivation of the pseudo-characteristic polynomial and the determination of Hopf bifurcation criteria, are employed to investigate the stability properties of steady-state solutions. The analysis reveals the critical role of multiple, dissimilar delays in modulating system dynamics and inducing bifurcations. Numerical simulations, conducted in conjunction with analytical results, reveal the emergence of periodic self-sustained oscillations and intermittent chaotic behaviour. These observations emphasise the intricate relationship between temporal heterogeneity and the stability landscape of systems exhibiting QS-inspired dynamics. This interplay highlights the capacity for temporal variations to induce complex dynamical transitions within such systems. These findings assist to the comprehension of temporal dynamics within these and related systems, and may contribute to the development of strategies aimed at modulating intercellular communication and engineering synthetic biological systems with temporal control.
@@ 1 Introduction
Quorum sensing (QS) is a fundamental mechanism of microbial communication that was first proposed in the early 1970s to explain the bioluminescence observed in marine bacteria, particularly Vibrio fischeri [29]. Over the decades, QS has emerged as a sophisticated strategy that enables bacterial populations to synchronise their behaviour and enhance collective fitness. This process allows bacteria to assess their local density by monitoring the concentration of signalling molecules known as autoinducers. Once the concentration of autoinducers surpasses a critical threshold, a coordinated response is initiated, conferring advantages such as enhanced virulence [21, 23] and biofilm formation [17].
@@
The study of QS mechanisms has been benefited from diverse mathematical modelling approaches, each offering unique insights into the underlying biological dynamics. These approaches span deterministic models, which describe average system behaviours; stochastic models, which account for intrinsic fluctuations; and hybrid models that combine both perspectives. For instance, Velázquez et al. [31] employed a continuous-time Markov process to investigate virulence in plant-pathogen interactions on leaf surfaces. Their model integrates linear birth rates, logistic death-migration processes, and an autocatalytic mechanism for acyl homoserine lactone autoinducers, revealing an inverse relationship between QS efficiency and the diffusion of autoinducers. Similarly, Frederick et al. [11] used a reaction-diffusion model with density-dependent diffusion coefficients to study extracellular polymeric substance (EPS) production in biofilms. They demonstrated how QS-regulated EPS production facilitates biofilms’ transition from a colonisation phase to a protective state, underscoring the adaptive benefits of QS in biofilm environments.
Delays and anomalous diffusion have also been key features in deterministic modelling approaches. Kuttler et al. [19] explored anomalous diffusion processes to model delayed substance transport caused by transcriptional delays. Chen et al. [5], Elowitz et al. [7], and Ojalvo et al. [12] studied QS pathways using gene regulatory networks with feedforward and negative feedback loops. Their models incorporated transcriptional time delays to explore phenomena such as Hopf bifurcations and sustained oscillations. These studies showed that transcriptional delays can act as critical parameters, leading to oscillatory and complex behaviours, particularly in the context of QS-regulated drug delivery systems for cancer therapy.
In bacterial QS, specialised receptor proteins detect autoinducers, triggering gene expression changes that orchestrate collective behaviours. These changes lead to a division into two sub-populations: motile (up-regulated), which engage in collective activities, and static (down-regulated), which remain inactive. However, transitions between these states are not instantaneous, as biochemical processes like transcription and translation introduce inherent time delays [22, 27, 38, 37]. These delays, which reflect processes such as signal production, diffusion, and cellular response, significantly influence the timing and robustness of QS-driven behaviours [28].
@@
The inclusion of time delays in QS models is crucial for accurately capturing the temporal dynamics of bacterial communication. For example, Barbarossa et al. [1] developed a delay differential equations (DDE) model for Pseudomonas putida to explain experimental observations of autoinducer production. Their results supported the hypothesis of an enzyme degrading autoinducers into an inactive form, thereby introducing a feedback delay. Time delays have been shown to induce oscillatory behaviours, as demonstrated in [4, 3], where the amplitude and frequency of oscillations depended on the delay length. Furthermore, bifurcation analyses of QS models have revealed transitions between stable steady states, periodic oscillations, chaotic behaviours (e.g. [24, 16, 18]), and multi-stability [35]. These features align with experimental observations where bacterial populations exhibit synchronised or desynchronised behaviours under varying environmental conditions [14].
@@
In the present manuscript, we analyse a nonlinear DDE quorum sensing system in which motile and static bacterial sub-populations respond differently to autoinducers due to distinct response times. This is characterised by the introduction of two independent delay parameters, $\tau_{1}>0$ and $\tau_{2}>0$, corresponding to the motile and static sub-populations, respectively. A local stability analysis is conducted by varying these delay parameters, identifying diverse bifurcations and the conditions under which they arise. These findings contribute to a deeper understanding of how time delays shape the dynamics of QS systems, particularly in the context of bifurcations and stability transitions.
@@
The manuscript is structured as follows: Section 2 introduces the nonlinear time-delayed QS system under investigation, detailing the model’s assumptions and highlighting the activator-inhibitor framework employed to represent the dynamic interplay between motile and static bacterial subpopulations. Section 3 rigorously analyzes the existence and local stability properties of the system’s steady states, deriving the pseudo-characteristic polynomial and formulating conditions for the emergence of purely imaginary eigenvalues, thereby facilitating the analysis of Hopf bifurcations. Section 4 presents the results of extensive numerical simulations, exploring the system’s dynamical behavior across a range of parameter values and illustrating the profound influence of varying delay parameters on the system’s dynamics. A detailed analysis of the observed intermittent chaotic behaviour, including a discussion of its characteristics and underlying mechanisms, is presented in Section 5. Finally, Section 6 concludes the manuscript with a comprehensive discussion of the key findings, emphasising their implications for advancing our understanding of QS-inspired systems and, in particular, the emergence of complex dynamical behaviour in the presence of multiple delays.
@@ 2 Nonlinear time delayed Quorum Sensing system
where $w$, $v$, and $u$ represent the concentrations of autoinducers, static bacteria, and motile bacteria subpopulations, respectively. The parameters are defined as follows: (i) $K$is the saturation parameter, modulating the regulatory effects of autoinducer concentrations, (ii) $a$and $\varepsilon a$ denote the production rates of autoinducers and the bacterial population, respectively, (iii) $b$is the decay rate of the bacterial population, and (iv) $c$and $d$ describe the decay and production rates of autoinducers.
@@
$a$ $b$ $\varepsilon$ $c$ $d$ $K$ $\tau_{1}$ $\tau_{2}$
$0.0428$ $[1.2,1.5]$ $0.1$ $0.16$ $0.0522$ $0.0435$ $[0,500]$ $[0,500]$
Table 1: Parameter values set.
As the bacterial response to stimuli communicated by autoinducers is not necessarily instantaneous, it is reasonable to hypothesise that this response may depend on intrinsic receptor properties within each sub-population or on bacterial activities carried out at earlier times.
To account for these temporal effects, the original system (1) is extended by incorporating delayed autoinducer concentrations into the growth rates. Specifically, let $w_{1}=w(t-\tau_{1})$ and $w_{2}=w(t-\tau_{2})$, where $\tau_{1}>0$ and $\tau_{2}>0$ represent the delay times associated with the non-instantaneous responses of motile and static bacteria, respectively. The modified QS system is then given by:
@@
Notice that the inclusion of delays $\tau_{1}$ and $\tau_{2}$ introduces additional complexity to the system dynamics, as these delays reflect the temporal lag between autoinducer production and its regulatory effects on bacterial populations. The delayed terms $w_{1}$ and $w_{2}$ capture the influence of past autoinducer concentrations on the current growth rates of static and motile bacteria, respectively. Our analysis focuses on examining the dynamic impact of these delays, particularly how they influence stability and potential bifurcation phenomena.
@@
Biologically meaningful steady-states of system (2) correspond to constant solutions that lie within the first octant. These steady-states can be determined by solving the following algebraic system:
@@
To refine this further, the Sturm Theorem [36] can be applied to determine the number of distinct real roots of (5) within the interval $|u|<\sqrt{c/d}$. Specifically, the difference in sign variations of the Sturm sequence at the interval’s endpoints provides the number of real roots in the interval. Using the parameter values given in Table 1 with $b=1.4$, the signs of the Sturm sequence are calculated at the ends of the interval of consideration as follows: (a) at $u=-1.75$, the sequence is $S(-1.75)=\{+,-,-,+,-,+,-\}$ , resulting in $V(-1.75)=5$ sign changes; (b) at $u=1.75$, the sequence is $S(1.75)=\{+,+,-,-,-,-,-\}$ , resulting in $V(1.75)=1$ sign change. Hence, the number of real roots within the interval is given by $V(-1.75)-V(1.75)=4$. Combining this result with Descartes’ rule of signs confirms that the roots of $f_{eq}(u)$ comprise three positive, one negative, and two complex conjugate roots. In consequence, there are three steady-states lying within the first octant.
@@
Now, to analyse the local stability properties of the system, let us denote the positive steady-states as $E_{1},E_{2},E_{3}\in\mathbb{R}^{3}$. The pseudo-characteristic polynomial at these steady-states is given by (see, for instance, [26]):
@@
where $\mathbf{I}_{3\times 3}$ is the identity matrix, and the matrices $\mathbf{J}_{0}$, $\mathbf{J}_{\tau_{1}}$, and $\mathbf{J}_{\tau_{2}}$ are defined as follows:
@@
The matrix $\mathbf{J}_{0}$ corresponds to the Jacobian evaluated at the steady-state in the absence of delays, while $\mathbf{J}_{\tau_{1}}$ and $\mathbf{J}_{\tau_{2}}$ capture the contributions from delayed interactions associated with $w_{1}=w(t-\tau_{1})$ and $w_{2}=w(t-\tau_{2})$, respectively. These delay terms reflect the biological reality that bacterial responses to autoinducers may not be instantaneous, depending on the activity of each subpopulation or autoinducers response characteristics.
@@
Observe that the determinant in (6) gives rise to the quasi-polynomial expression
@@
here, $\boldsymbol{\tau}=(\tau_{1},\tau_{2})^{T}$ represents the delay vector, while $\mathbf{p}=(a,b,c,d,\varepsilon,K)^{T}$ encapsulates the system parameter vector.
@@
Condition (10a) ensures that $|\xi(\omega)|\leq 1$, which constitutes a necessary criterion for the solvability of equation (10b). Consequently, the roots of (8) correspond to values of $h\in\mathbb{R}$ and $\omega>0$ at which the functions $\xi(\omega)$ and $\cos(\omega h-\alpha(\omega))$ intersect. This behaviour is depicted in Figure 1(a), which illustrates the intersection points for the steady-state $E_{1}$ at three distinct values of $h$. A similar pattern is observed for the steady-state $E_{2}$, as is shown in Figure 1(b). Notably, the orders of magnitude of $h$ at which these intersections occur differ significantly between the two steady-states: for $E_{1}$, intersections arise at values of $h$ on the order of $10^{2}$, whereas for $E_{2}$, they occur at values two orders of magnitude greater. In both cases, the frequency of oscillations in $\cos(\omega h-\alpha(\omega))$ increases proportionally with $h$, facilitating the number of intersections, which suggests that the emergence of purely imaginary roots is strongly influenced by the introduction of two distinct delay terms, particularly when their magnitudes differ considerably. On the other hand, the steady-state $E_{3}$ is evaluated in the functions $\xi(\omega)$ and $\cos(\cdot)$. In this case, it is disclosed that $|\xi(\omega)|>1$ for all $\omega\geq 0$. Consequently, condition (10a) is hence not satisfied, preventing the existence of intersection solutions for (10b), which results in confirming that $E_{3}$ remains stable for all $\tau_{1},\tau_{2}>0$.
@@
Figure 1: Graphs of $\xi(\omega)$ (black) and $\cos(\cdot)$ (blue), along with their intersection points, are shown for the steady-states: (a) $E_{1}$ at $h=124,262,400$; (b) $E_{2}$ at $h=2\times 10^{4},3\times 10^{4},4\times 10^{4}$. All other parameter values are as in Table 1.

### v3

@@ Abstract
This study analyses the dynamical consequences of heterogeneous temporal delays within a quorum sensing-inspired (QS-inspired) system, specifically addressing the differential response kinetics of two sub-populations to signalling molecules. A nonlinear delay differential equation (DDE) model, predicated upon an activator-inhibitor framework, is formulated to represent the interspecies interactions. Key analytical techniques, including the derivation of the pseudo-characteristic polynomial and the determination of Hopf bifurcation criteria, are employed to investigate the stability properties of steady-state solutions. The analysis reveals the critical role of multiple, dissimilar delays in modulating system dynamics and inducing bifurcations. Numerical simulations, conducted in conjunction with analytical results, reveal the emergence of periodic self-sustained oscillations and intermittent chaotic behaviour. These observations emphasise the intricate relationship between temporal heterogeneity and the stability landscape of systems exhibiting QS-inspired dynamics. This interplay highlights the capacity for temporal variations to induce complex dynamical transitions within such systems. These findings assist to the comprehension of temporal dynamics within these and related systems, and may contribute to the development of strategies aimed at modulating intercellular communication and engineering synthetic biological systems with temporal control.
@@ 1 Introduction
Quorum sensing (QS) is a fundamental mechanism of microbial communication that was first proposed in the early 1970s to explain the bioluminescence observed in marine bacteria, particularly Vibrio fischeri [41]. Over the decades, QS has emerged as a sophisticated strategy that enables bacterial populations to synchronise their behaviour and enhance collective fitness. This process allows bacteria to assess their local density by monitoring the concentration of signalling molecules known as autoinducers. Once the concentration of autoinducers surpasses a critical threshold, a coordinated response is initiated, conferring advantages such as enhanced virulence [28, 30], where QS triggers the expression of virulence factors, enabling coordinated attack on a host once a sufficient bacterial community is established, as seen in Pseudomonas aeruginosa infections [37, 39]. Another advantage is the colony capacity to reach biofilm formation [24], a structured community of bacteria encased in a self-produced matrix, which enhances their resistance to antibiotics and host immune responses, exemplified by Staphylococcus aureus biofilms in chronic infections [44, 46]. That is, understanding and targeting QS pathways presents promising avenues for novel antimicrobial strategies that aim to disrupt bacterial cooperation rather than directly killing them, potentially mitigating the rise of antibiotic resistance.
@@
The study of QS mechanisms has been benefited from diverse mathematical modelling approaches, each offering unique insights into the underlying biological dynamics. These approaches span deterministic models, which describe average system behaviours; stochastic models, which account for intrinsic fluctuations; and hybrid models that combine both perspectives. For instance, Gorochowski [19] examines the broader scope of agent-based models in synthetic biology, including their benefits and challenges. Meanwhile, Velázquez et al. [47] employed a continuous-time Markov process to specifically investigate virulence in plant-pathogen interactions on leaf surfaces; their detailed model incorporated linear birth rates, logistic death-migration processes, and an autocatalytic mechanism for acyl homoserine lactone autoinducers, demonstrating an inverse relationship between QS efficiency and autoinducer diffusion. Similarly, Frederick et al. [16] used a reaction-diffusion model with density-dependent diffusion coefficients to study extracellular polymeric substance (EPS) production in biofilms. They demonstrated how QS-regulated EPS production facilitates biofilms’ transition from a colonisation phase to a protective state, underscoring the adaptive benefits of QS in biofilm environments.
In particular, deterministic modelling has advanced our understanding of biological systems by incorporating the critical roles of temporal delays among other transport features not under the scope of this manuscript. Specifically, the inherent delays associated with intracellular processes, such as transcription and translation, can profoundly influence the dynamic behaviour of regulatory networks.
The particular case of transcriptional time delays has been a subject of intense investigation. The intricate gene regulatory networks governing such mechanism, often featuring feedforward and negative feedback motifs, are particularly susceptible to the influence of these delays. Seminal works by Chen et al. [8], Elowitz et al. [12], and Ojalvo et al. [17] demonstrated how transcriptional delays may act as crucial parameters that affect the stability and oscillatory properties of QS circuits. Their models revealed that specific ranges of time delays can induce dynamical events such as Hopf bifurcations, leading to the emergence of sustained oscillations in the concentrations of signalling molecules and downstream effectors. This temporal orchestration has direct implications for the functional outcomes of QS, such as the precisely timed release of virulence factors in pathogenic bacteria or the coordinated formation and dispersal of biofilms.
@@
The inclusion of time delays in QS models is crucial for accurately capturing certain key temporal dynamics of bacterial communication. For example, Barbarossa et al. [4] developed a delay differential equations (DDE) model for Pseudomonas putida to explain experimental observations of autoinducer production. Their results supported the hypothesis of an enzyme degrading autoinducers into an inactive form, thereby introducing a feedback delay. Time delays have been shown to induce oscillatory behaviours, as demonstrated in [7, 6], where the amplitude and frequency of oscillations depended on the delay length. Furthermore, bifurcation analyses of QS models have revealed transitions between stable steady-states, periodic oscillations, chaotic behaviours (e.g. [32, 23, 31]), and multi-stability [57]. These features align with experimental observations where bacterial populations exhibit synchronised or desynchronised behaviours under varying environmental conditions [20].
Moreover, comprehension of delay-induced dynamics in QS has spurred innovation in biomedical engineering. The ability to harness QS-regulated systems for applications like controlled drug delivery in cancer therapy hinges on the precise temporal control afforded by these inherent biological delays. As demonstrated by You et al. [58], cell-cell communication mechanisms, such as QS, can be engineered to achieve programmed population control, highlighting the potential for temporally regulated biological processes. Building upon this, Tamsir et al. [53] showcased robust multicellular computing using synthetic biology, where QS enables coordinated behaviour within a cell population, demonstrating the capacity for controlled actions in biological systems. The inherent time lags in QS responses are crucial for achieving this level of control.
@@
In bacterial QS, certain receptor proteins perceive autoinducers, initiating intricate gene expression programmes that govern collective behaviours. This often results in the emergence of functionally distinct sub-populations: a motile or active fraction, where genes related to collective activities are significantly up-regulated, and a static or reservoir fraction, where these genes are correspondingly down-regulated. The motile subpopulation actively contributes to processes like biofilm formation, virulence factor secretion, or bioluminescence, while the static subpopulation may serve as a reserve, potentially contributing to population survival under stress or acting as a source for future active cells. However, the transitions between these dynamic states are not immediate consequences of autoinducer detection. Instead, inherent time delays, arising from fundamental biochemical processes such as the transcription of genes, the translation of messenger RNA into proteins, and the maturation or activation of these proteins, introduce a temporal dimension to QS responses [29, 38, 61, 60, 40]. These delays, which encapsulate the time scales of signal production, extracellular diffusion, receptor binding, intracellular signalling cascades, and the eventual transcriptional and translational responses, critically modulate the timing, synchronisation, and robustness of QS-mediated collective behaviours [40]. Even though there is intense research of time delay systems occurring in other fields as reported in [2, 3], for instance, we focus on understanding these delay-dependent dynamics as crucial for deciphering the complex temporal organisation particularly within bacterial communities and for developing strategies to manipulate QS-controlled processes.
@@
In the present manuscript, we analyse a nonlinear DDE-QS system wherein motile and static bacterial sub-populations exhibit differential responses to autoinducers due to distinct response times. This is characterised by the introduction of two independent delay parameters, $\tau_{1}>0$ and $\tau_{2}>0$, corresponding to the motile and static sub-populations, respectively. A local stability analysis is conducted by varying these delay parameters, identifying diverse bifurcations and the conditions under which they arise. These findings contribute to the understanding of how time delays shape the dynamics of QS systems, particularly concerning bifurcations and stability transitions. Furthermore, the emergence of chaotic dynamics from such transitions holds the potential to significantly clarify crucial bacterial population behaviour, potentially enabling the identification of key conditions leading to disruptive QS, among other phenomena. This parallels how chaos theory has illuminated key features in other systems, as explored in [9, 45].
The theoretical framework for deterministic chaos in biological systems characterised by nonlinear feedback and time delays was established by the seminal works of Mackey and Glass [32] and Guevara et al. [23]. These investigations demonstrated that complex temporal dynamics could arise endogenously from such system architectures, a concept further explored in the context of biological oscillators [35].
Mathematical modelling, as evidenced in studies demonstrating bifurcations and chaos in synthetic QS systems and analysing chaotic behaviours in delayed reaction-diffusion QS models, reveals the potential for chaotic regimes within QS networks under specific parametric conditions. This is consistent with the broader understanding that nonlinearities and time delays inherent in QS mechanisms can lead to complex dynamics [25, 26]. Complementing these findings, Harris et al. [31] provide a mechanistic understanding of these complex behaviours by detailing the role of local and global bifurcations in facilitating transitions between stable states and potentially chaotic invariant sets. Their analysis of bifurcation phenomena elucidates how parameter variations within QS systems can lead to dynamic instability and complex outputs.
The broader functional implications of such dynamic complexity for bacterial adaptation and resilience are reflected in the investigation of hyperchaos in coupled QS systems. Stankevich and Volkov [50] suggest that interactions between bacterial units or QS circuits can amplify dynamical complexity. The potential for chaotic dynamics to offer a form of robustness or adaptability for bacterial communities, allowing for a wider exploration of behavioural states, aligns with the theoretical possibilities discussed, for instance, in [6, 25, 26].
@@
The manuscript is structured as follows: Section 2 introduces the nonlinear time-delayed QS system under investigation, detailing the model’s assumptions and highlighting the activator-inhibitor framework employed to represent the dynamic interplay between motile and static bacterial sub-populations. Section 3 rigorously analyses the existence and local stability properties of the system’s steady-states, deriving the pseudo-characteristic polynomial and formulating conditions for the emergence of purely imaginary eigenvalues, thereby facilitating the analysis of Hopf bifurcations. Section 4 presents the results of extensive numerical simulations, exploring the system’s dynamical behaviour across a range of parameter values and illustrating the profound influence of varying delay parameters on the system’s dynamics. A detailed analysis of the observed intermittent chaotic behaviour, including a discussion of its characteristics and underlying mechanisms, is presented in Section 5. Finally, Section 6 concludes the manuscript with a comprehensive discussion of the key findings, emphasising their implications for advancing our understanding of QS-inspired systems and, in particular, the emergence of complex dynamical behaviour in the presence of multiple delays.
@@ 2 Nonlinear time delayed Quorum Sensing system
where $w$, $v$, and $u$ represent the concentrations of autoinducers, static bacteria, and motile bacteria sub-populations, respectively. The parameters are defined as follows: (i) $K$is the saturation parameter, modulating the regulatory effects of autoinducer concentrations; (ii) $a$and $\varepsilon a$ denote the production rates of autoinducers and the bacterial population, respectively; (iii) $b$is the decay rate of the bacterial population; and (iv) $c$and $d$ describe the decay and production rates of autoinducers.
@@
Given that bacterial responses to autoinducer-mediated stimuli are not immediate, it is plausible that the ensuing cellular activities are contingent upon intrinsic receptor characteristics within distinct sub-populations or on physiological processes enacted at preceding times. While acknowledging the biological limitations of relying solely on constant delays, their initial and selective use in QS models is supported by some pragmatic advantages. Firstly, the mathematical simplicity and analytical tractability of DDEs with constant delays provide a robust framework for initial investigations into system dynamics, which allows to readily explore stability landscapes, identify potential bifurcations, and characterise oscillatory regimes, offering qualitative insights into the fundamental impact of temporal lags on QS behaviour. Secondly, employing constant delays can serve as a crucial first-order approximation of the system’s temporal characteristics. Upon capturing the non-instantaneous nature of signal transduction and gene expression, this simplified model provides a foundational understanding of how time scales influence QS responses, particularly when detailed kinetic data for more complex delay formulations are unavailable. Finally, in specific contexts when the modelling focus is on emergent macroscopic population-level phenomena as is in the present case, constant delays can offer a reasonably accurate representation without the added complexities of distributed or state-dependent delays. Consequently, while not offering a comprehensive depiction of biological reality, constant delays can serve as an entry point for exploring the temporal intricacies of QS.
To account for these temporal effects as well as the delays features already described in Section 1, the original system (1) is extended by incorporating delayed autoinducer concentrations into the growth rates. Specifically, let $w_{1}=w(t-\tau_{1})$ and $w_{2}=w(t-\tau_{2})$, where $\tau_{1}>0$ and $\tau_{2}>0$ represent the delay times associated with the non-instantaneous responses of motile and static bacteria, respectively. The modified QS system is then given by:
@@
Notice that the inclusion of delays $\tau_{1}$ and $\tau_{2}$ introduces additional complexity to the system dynamics, as these delays reflect the temporal lag between autoinducer production and its regulatory effects on bacterial populations. The delayed terms $w_{1}=w(t-\tau_{1})$ and $w_{2}=w(t-\tau_{2})$ capture the influence of past autoinducer concentrations on the current growth rates of static and motile bacteria, respectively. Our analysis focuses on examining the dynamic impact of these delays, particularly how they influence stability and potential bifurcation phenomena.
$a$ $b$ $\varepsilon$ $c$ $d$ $K$ $\tau_{1}$ $\tau_{2}$
$0.0428$ $[1.2,1.5]$ $0.1$ $0.16$ $0.0522$ $0.0435$ $[0,500]$ ${[0,1000]}$
Table 1: Parameter values set as given in [31], and a wide range for delay parameters $\tau_{1,2}$.
@@
Biologically meaningful steady-states of system (2) correspond to constant solutions that lie within the first octant. Necessarily, such a solutions are determined by the condition where the time derivatives on the left-hand side of either system (1) or system (2) vanish. That is, these steady-states are thus determined by solving the following algebraic system:
@@
To refine this further, the Sturm Theorem can be applied to determine the number of distinct real roots of (5) within the interval $|u|<\sqrt{c/d}$. Using the parameter values given in Table 1 with $b=1.4$, the signs of the Sturm sequence are calculated at the ends of the interval of consideration. Consequently, the number of real roots within the interval is four. Combining this result with Descartes’ rule of signs confirms that the roots of $f_{eq}(u)$ comprise three positive, one negative, and two complex conjugate roots. In consequence, there are three steady-states lying within the first octant. This yields to the following proposition
Proposition 3.1.
For system (3), its associated monic polynomial of degree six, defined in (5), possesses three positive real roots, one negative real root, and two complex conjugate roots if and only if $(a+1)d>ba\varepsilon cK$. Under this condition, and within the interval $|u|<\sqrt{c/d}$, all positive steady-states of system (2) are determined.
@@
Results presented in [31] reveal oscillatory dynamics within a non-delayed framework in the context of this model, but significantly, they also uncover chaotic dynamics explained by homoclinic Shilnikov bifurcations (see, for instance, [1]), which arise from a primary Hopf bifurcation. These bifurcation findings are shown to be dependent on external rather than internal parameters.
In contrast, our approach here focuses on disclosing the consequences of slowly varying internal delay parameters. Unlike the simplification of a single, constant delay—which can introduce temporal effects and oscillations absent in non-delayed systems—we acknowledge that real biological systems often exhibit a distribution of delays. Consequently, a single constant value inherently fails to capture the richer information encoded within these more realistic delay distributions. Such a simplification can prevent the model from reproducing more complex dynamics, including intricate oscillatory patterns, bifurcations triggered by delay magnitude, or multi-stability, which can arise when a range of delays is considered. Thus, our initial exploration using a single delay (not shown) proved inadequate to sufficiently portray the complex temporal regulation within the QS system under study.
That is, while Harris et al. in [31] provide a detailed analysis of Hopf bifurcations in the non-delayed QS model version, the general principles are fundamental to understanding the more complex dynamics of DDEs. The presence of delays in QS systems adds a layer of complexity, making the study of bifurcations crucial for deciphering and potentially controlling bacterial behaviour. Let us recall the crucial points of Hopf bifurcation [36, 43, 52, 55]: (i) As can be seen further, the stability of a steady-state is determined by a characteristic equation that is transcendental, unlike the polynomial equation in ordinary differential equations (ODEs). This transcendental equation has infinitely many roots (eigenvalues), making the analysis more intricate. (ii) Similar to ODEs, a Hopf bifurcation in DDEs occurs when a pair of complex conjugate eigenvalues of the linearised system cross the imaginary axis as a parameter slowly varies, which is known as a codimension-one bifurcation. However, in DDEs, this crossing can happen multiple times for different eigenvalue pairs as the delay is increased, leading to multiple secondary bifurcations and complex dynamics. (iii) The delay itself is a critical bifurcation parameter in DDEs. Varying the delay can induce oscillations and instability even if the steady-state for the corresponding ODE system is stable, as well as affecting the frequency of the emerging oscillations, as it is determined by the imaginary part of the eigenvalues crossing the imaginary axis. This is a key difference from ODEs, where the delay is absent. (iv) The stability of the bifurcating periodic solutions is determined by the direction of the eigenvalue crossing and the non-linear terms of the system. This can be analysed using normal form theory and the computation of Lyapunov coefficients, as shown, for instance, in [31] for the non-delayed case. We here propose an alternative approach.
Hence, to analyse the local stability properties of the system, let us denote the positive steady-states discovered in Section 3.1 as $E_{1},E_{2},E_{3}\in\mathbb{R}^{3}$.
We consider the nonlinear DDE system defined by $\dot{\mathbf{x}}=\mathbf{F}(\mathbf{x},\mathbf{x}_{\tau_{1}},\mathbf{x}_{\tau_{2}})$. Here, the non-delayed state vector is $\mathbf{x}=(w,v,u)\in\mathbb{R}^{3}_{+}$, and $\mathbf{x}_{\tau_{1}}=(w_{1},0,0)$ and $\mathbf{x}_{\tau_{2}}=(w_{2},0,0)$ denote the two time-delayed states, associated with positive time delays $\tau_{1},\tau_{2}>0$. The function $\mathbf{F}=(f_{1},f_{2},f_{3})$ is a differentiable vector-field, and its components correspond to the right-hand side of system (2). Analogous to the non-delayed case, a linearisation approach is applied to this retarded nonlinear system. This technique relies on constructing local estimations of the nonlinear states in the vicinity of steady-states. Let
$\left(\mathbf{J}_{0}\right)_{i,j}=\dfrac{\partial f_{i}}{\partial x_{j}}\,,\quad i,j=1,2,3\,,$
denote the matrix of system partial derivatives with respect to the non-delayed states. Similarly, let matrices
$(\mathbf{J}_{\tau_{k}})_{i,j}=\dfrac{\partial f_{i}}{\partial x_{\tau_{kj}}}\,,\quad i,j=1,2,3\quad\textrm{and}\quad k=1,2\,,$
account for the local contributions of the non-immediate variables arising from the time delays $\tau_{1},\tau_{2}>0$. The ansatz of an exponential solution for the linearised system necessitates a nontrivial evolution near steady-states only when there exist complex values $\lambda$ such that the so-called pseudo-characteristic polynomial satisfies (see [36]):
@@
where $\mathbf{I}_{3\times 3}$ is the identity matrix and exponential factors arise given the dependence on the history of variables within past intervals defined by $\tau_{1}$ and $\tau_{2}$.
In the particular case of QS time delayed system (2), matrices $\mathbf{J}_{0}$, $\mathbf{J}_{\tau_{1}}$, and $\mathbf{J}_{\tau_{2}}$ are given by
@@
The matrix $\mathbf{J}_{0}$ corresponds to the Jacobian evaluated at the steady-state in the absence of delays, while $\mathbf{J}_{\tau_{1}}$ and $\mathbf{J}_{\tau_{2}}$ capture the contributions from delayed interactions associated with $w_{1}$ and $w_{2}$, respectively. These delay terms reflect the biological reality that bacterial responses to autoinducers may not be instantaneous, depending on the activity of each subpopulation or autoinducers response characteristics.
@@
Observe that, upon substituting (6b) and (6c) into (6a), we get the pseudo-polynomial
@@
here, $\boldsymbol{\tau}=(\tau_{1},\tau_{2})$ represents the delay vector, while $\mathbf{p}=(a,b,c,d,\varepsilon,K)$ encapsulates the system parameter vector.
@@
As a direct consequence of this result, we state:
Proposition 3.2.
The pseudo-polynomial defined by equation (7) possesses purely imaginary roots if and only if conditions (10) are met.
@@
Figure 1: Existence of purely imaginary eigenvalues for the pseudo-polynomial (7). Such solutions are guaranteed for $\omega>0$ values that satisfy (10); that is, where functions $\xi(\omega)$ (black curve) and $\cos(\omega h-\alpha(\omega))$ with fixed $h$ (blue curve) intersect. Cases for equilibria $E_{1}$ and $E_{2}$ are considered separately. (a) Analysis of intersections at $E_{1}$ for $h$ values of $124,262,400$. (b) Similar solution visualisations for $E_{2}$ with $h=2\times 10^{4},3\times 10^{4}.4\times 10^{4}$. Parameter values as in Table 1.
Condition (10a) ensures that $|\xi(\omega)|\leq 1$, which constitutes a necessary criterion for the solvability of equation (10b). Consequently, the roots of (8) correspond to values of $h\in\mathbb{R}$ and $\omega>0$ at which the functions $\xi(\omega)$ and $\cos(\omega h-\alpha(\omega))$ intersect. This behaviour is depicted in Figure 1(a), which illustrates the intersection points for the steady-state $E_{1}$ at three distinct values of $h$. A similar pattern is observed for the steady-state $E_{2}$, as is shown in Figure 1(b). Notably, the orders of magnitude of $h$ at which these intersections occur differ significantly between the two steady-states: for $E_{1}$, intersections arise at values of $h$ on the order of $10^{2}$, whereas for $E_{2}$, they occur at values two orders of magnitude greater. In both cases, the frequency of oscillations in $\cos(\omega h-\alpha(\omega))$ increases proportionally with $h$, facilitating the number of intersections, which suggests that the emergence of purely imaginary roots is strongly influenced by the introduction of two distinct delay terms, particularly when their magnitudes differ considerably. This suggests a greater robustness in the stability of equilibrium $E_{2}$ compared to $E_{1}$. On the other hand, the steady-state $E_{3}$ is evaluated in the functions $\xi(\omega)$ and $\cos(\cdot)$. In this case, it is disclosed that $|\xi(\omega)|>1$ for all $\omega\geq 0$. Consequently, condition (10a) is hence not satisfied, preventing the existence of intersection solutions for (10b), which results in confirming that $E_{3}$ remains stable for all $\tau_{1},\tau_{2}>0$.